OpenStudy (anonymous):

Will somebody help me and explain how to find the rate of change of 5x+3y=-2......PLEASE PLEASE PLEASE

OpenStudy (anonymous):

Well, turn it into a y = mx + b equation first.

OpenStudy (anonymous):

3y = -5x - 2 y = -5/3x - 2/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so in the second equation, whih one is the slope?

OpenStudy (anonymous):

-5/3?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so the slope is the same thing as the rate of change. so there's your answer :)

OpenStudy (anonymous):

O MY GOSH THANK YOU SO MUCH.... but one more question?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

on graphs would the rate of change be the slope as well?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

using a graph, if you don't know the slope formula, use this: slope is rise over run. so on the graph of the line, find an point where both x and y values are integers. Take a look at the rise (number of units up) and the run(number of horizontal units) to another integer point. tthe rise over the run would be the slope.

OpenStudy (anonymous):

and you're welome, anytime :)

OpenStudy (anonymous):

*welcome